According to a recent survey, the salaries of assistive professors have a mean of $41,750 and a standard deviation of $7500. Assuming that the salaries of assistive assessors follow a normal distribution, find the proportion of assistance professors who are at most $48,000. Where are you answer to at least four decimal places.
1
Expert's answer
2020-12-22T18:54:39-0500

μ = 41750

x = 48000

σ = 7500

"P(x≤48000) = P(Z \leq \frac{x-μ}{σ}) \\

= P(Z \leq \frac{41750-48000}{7500}) \\

=P(Z \leq \frac{-6250}{7500}) \\

= P(Z\leq-0.8333) \\

= 0.2024"